# What Foundation Models can Bring for Robot Learning in Manipulation : A Survey

### v6

@@ 1 Introduction
Researchers aim to create universal robots that can seamlessly integrate into human life to boost productivity, much like those depicted in the movie ‘I, Robot’. However, a key hurdle in achieving this lies in the robots’ ability to manipulate objects in their unstructured environments according to different tasks. There is abundant literature available for improving the general manipulation capability of robots, which can be roughly categorized into model-based and learning-based approaches (Zarrin et al. (2023)). The real world is too diverse for universal robots and they must adapt to unstructured environments and arbitrary objects to manipulate effectively. Therefore, learning-based methods are crucial for manipulation tasks (Kleeberger et al. (2020)).
The predominant methodologies in learning-based approaches are deep learning, reinforcement learning and imitation learning. Learning-based methods have spanned from acquiring specific manipulation skills through labeled datasets like human demonstration, to acquiring abstract representations of manipulation tasks conducive to high-level planning, to exploring an object’s functionalities through interaction and encompassing various objectives in between (Kroemer et al. (2021)). However, challenges persist, including 1) unnatural interaction with humans; 2) high-cost data collection; 3) limited perceptual capability; 4) non-intelligent hierarchy of skills; 5) inaccurate pre- and post-conditions & post-hoc correction; 6) unreliable skill learning; 7) poor environment transition (Hu et al. (2023b)).
Foundation models are primarily pretrained on vast internet-scale datasets, enabling them to be fine-tuned for diverse tasks. Their significant advancements in vision and language processing contribute to mitigating the aforementioned challenges. Based on Firoozi et al. (2023) and considering the different input modalities and functionalities of the models, we categorize foundation models into the following six types.
@@
    Large Language Models (LLMs) like BERT (Devlin et al. (2018)), GPT-3 (Brown et al. (2020)) demonstrate the capability to generate coherent chains of thought.
@@
    Visual Foundation Models (VFMs) like SAM (Kirillov et al. (2023)) demonstrate strong segmentation capability for open-set objects.
@@
    Visual Generative Models (VGMs) like DALL-E (Ramesh et al. (2021)), Zero-1-to-3 (Liu et al. (2023e)) and Sora (Brooks et al. (2024)), demonstrate the capability to generate 2D images, videos or 3D meshes through text or images.
@@
    Visual-Language Models (VLMs) like GPT-4V (Achiam et al. (2023)), CLIP (Radford et al. (2021)) showcase robust comprehension of both vision and language, such as open-set image classification and visual question answering.
@@
    Large Multimodal Models (LMMs) expand their scope beyond vision and language to create novel categories of foundation models incorporating additional modalities, such as ULIP (Xue et al. (2023a)) aligns point cloud representation to the pre-aligned image-text feature space. VLMs are a type of LMMs (Firoozi et al. (2023)). However, due to the current literature focusing more on VLMs, this paper will treat VLMs as a separate category. To avoid confusion, LMMs in this paper refer to those that include images, language, and more modalities.
@@
    Robotic-specific Foundation Models (RFMs), like RT-X (Padalkar et al. (2023a)). Internet-scale dataset, such as images and text data, are suitable for pre-training visual and language models, but lack task level manipulation data. Therefore, researchers aim to train end-to-end RFMs by collecting task-level manipulation datasets to enable observations-to-action mapping.
@@
These findings underscore the potential of embedding foundation models into manipulation tasks as a viable path toward achieving general manipulation capability. However, we do not believe that a single foundation model alone can achieve general manipulation capability. Although RFMs currently represent a single-model end-to-end training approach, ensuring safety and stability, particularly in achieving an over 99% success rate in manipulation tasks, remains a challenge. Achieving over a 99% success rate in manipulation tasks is crucial, as human manipulation success rates are around 99%. Without this level of accuracy, robots can’t replace humans (Kumar (2023)). Therefore, drawing inspiration from the development of autonomous driving systems (Hu et al. (2023c)), achieving general manipulation capability necessitates an overarching framework that encompasses multiple functional modules, with different foundation models assuming distinct roles in facilitating general manipulation capability.
@@
The ultimate general manipulation framework should be able to interact with human or other agent and control whole-body to manipulate arbitrary objects in open-world scenarios and achieve diverse manipulation tasks (McCarthy et al. (2024)). Drawing from Kroemer et al. (2021) and this general manipulation definition, we propose a comprehensive framework for general manipulation. However, the interaction between robot and human involves not only recognizing intentions but also learning new skills or improving old skills from human experts in the external world. Open-world scenarios may be static or dynamic. Objects can be either rigid or deformable. Task objectives can vary from short-term to long-term. Furthermore, tasks may necessitate different degrees of precision with respect to contact points and applied forces/torques. Although there are many challenges, achieving general manipulation can be accomplished through multiple stages. We designate the restriction of the robot’s learning capability to improving old skills and to manipulating rigid objects in static scenes in order to achieve short-horizon task objectives with low precision requirements for contact points and forces/torques as Level 0 (L0). At the same time, we believe that improving the algorithm performance of different modules in the framework can support the transition from the L0 stage to the final general manipulation. Hence, we aim to use this survey not only to enlighten scholars on the issues that foundation models can address in robot learning for manipulation but also to stimulate their exploration of the general manipulation framework and the role various foundation models can play in the general manipulation framework.
@@
Di Palo et al. (2023) and Firoozi et al. (2023) provide detailed descriptions of the application of foundation models in navigation and manipulation, but these lack thoughtful consideration of the relationship between foundation models across different applications. The survey most closely related to this paper is Xiao et al. (2023). Compared to this survey, our survey focuses on the contributions of foundation models to robot learning for manipulation, proposing a comprehensive framework and detailing how foundation models can address challenges in each module of the framework.
@@ 2 Framework of Robot Learning for General Manipulation
Over the past decade, there has been a significant expansion in research concerning robot manipulation, with a focus on leveraging the growing accessibility of cost-effective robot arms and grippers to enable robots to interact directly with the environment in pursuit of their objectives. As the real world encompasses extensive variation, a robot cannot expect to possess an accurate model of its unstructured environment, the objects within it, or the skills necessary for manipulation in advance (Kroemer et al. (2021)).
Early stage, robot manipulation is defined as learning a policy $\Pi$ through deep learning, reinforcement learning, or imitation learning, etc. This policy controls the robot’s joint movements and executes tasks based on observations of the environment and the robot’s state $S$, mapping to actions $\alpha$. such as Rlafford (Geng et al. (2023b)) and Graspnet (Fang et al. (2020b)) take point cloud as input and output the target pose. This process is represented by the Skill Execution module, as shown in Fig. 2.
In the mid-term, many tasks in robotics require a series of correct actions, which are often long-horizon tasks. For example, making a cup of tea with a robot involves multiple sequential steps such as boiling water, adding a tea bag, pouring hot water, etc. Learning to plan for long-horizon tasks is a central challenge in episodic learning problems (Wang et al. (2020b)). Decomposing tasks has several advantages. It makes learning individual skills more efficient by breaking them into shorter-horizon, thus aiding exploration. Reusing skills in multiple settings can speed up learning by avoiding the need to relearn elements from scratch each time. Researchers train a hierarchy model to decompose the task into a sequence of subgoals (Ahn et al. (2022)), and observe pre- and post-conditions to ensure that the prerequisites and outcomes of each subgoals are met (Cui et al. (2022)). These three processes are represented as the Hierarchy of Skills module $H$, the Pre-conditions Detection module $P$, and the Post-conditions Detection module $P$ in Fig. 2. However, detecting only task success with post-conditions detection is insufficient. It should also identify the reasons for task failure to help the robot self-correct and improve success rates. Therefore, we add a Post-hoc Correction module, as shown in Fig. 2.
Recently, researchers have realized that training policies require real-world interaction between the robot and environments, which inevitably increases the probability of unforeseen hazardous situations. Therefore, researchers aim to train the environment’s transition model $T$. Once the model is fitted, robot can generate samples based on it, significantly reducing the frequency of direct interaction between the robot and environments (Liu et al. (2024e)). This process is represented as the Transition module $T$ in Fig. 2.
@@
Figure 2: Framework of Robot Learning for General Manipulation. The Pre-conditions Detection module $P$ perceives the environment to identify objects and the affordances objects support. The Interaction module II receives instruction from a human or other agent. It uses perception information from the Pre-conditions Detection module $P$ to check for ambiguities in the instruction. If there are any ambiguities, it generates a question to clarify the instruction by asking the human or other agent. The Hierarchy of Skills module $H$ generates subgoals by using precise instruction from the Interaction module $I$ and perception information from the Pre-conditions Detection module $P$. Each subgoal is then passed to the Skill Execution module. In the Skill Execution module, Policy module $\Pi$ generates Action $\alpha$ based on the State $S$. To obtain the next state after executing the current action, State $S$ can either perceive it from the environment or use the Transition module $T$. To train the Skill Execution module, including the State module $S$, the Policy module $\Pi$ and the Transition module $T$, the Manipulation Data Generation module is required. This module provides a task-level manipulation dataset. When issues arise during execution, corrective instruction is sent to the Policy module $\Pi$ for manual adjustment. Policy module $\Pi$ modifies the current action to corrective action and saves corrective demonstration to the dataset for self-improvement of Policy module $\Pi$. After skill execution, Post-conditions Detection module $P$ determines the success of execution. If successful, proceed to the next subgoal; if not, the failure reason is conveyed to Post-hoc Correction module for self-correction.
@@
    Hierarchy of skills. This module takes as input the perception information about objects in the environment and their affordances for the task from the Pre-conditions Detection module $P$, as well as the precise instruction from the Interaction module $I$. It then produces a sequence of subgoals as output. The concept of ‘Hierarchy of skills’ often involves creating a sequence of subgoals (Song et al. (2023)). Each subgoal necessitates a skill, which may consist of one or multiple primitive actions (Zhang et al. (2023c)). For instance, tasks like filling the kettle with water, heating the water, and getting the tea leaves are examples of subgoals that robot needs to achieve in a specific order to fulfill the final goal as instructed.
@@
    State. The input to the State module is the current environment, objects and robot observation. States require the use of multiple sensors for perception. The output is the features of states. The states consists of robot proprioception $S_{robot}$, environment state $S_{e}$, and objects states $S_{o}$. The difference between $S_{e}$ and $S_{o}$ is analogous to the foreground and background of an image. $S_{robot}$ generally relates to the mechanical structure of the robot. Currently, there are limited studies focusing on the improvement of robot mechanical structure using foundation models, with Stella et al. (2023) being one of them. However, researches in this direction are scarce and still in their initial stages.
@@
    Transition. The Transition module $T$ takes an action generated by the Policy module $P$ as input. It outputs the next state after executing this action, thus helping to reduce the interaction between the robot and the real environment. UniSim (Yang et al. (2023b)) introduces the action-in-video-out framework as an observation prediction model. It takes the current action as input and produces the subsequent observation as output.
@@ 3 Human/Agent Interaction
There are two ways for human or other agent to interact with robot: 1) Providing task instruction to the robot to help it understand the task objective and complete the task independently (Khan et al. (2023)). 2) Collaborating with the human or other agent to complete tasks, sharing workspace information, and conveying corrective instruction when useful or error-correcting information is identified to optimize the robot’s current action (Lynch et al. (2023)).
@@
In addressing instruction ambiguity and text generation and comprehension challenges, SeeAsk (Mo et al. (2023)) utilizes CLIP’s perceptual module to identify objects in the scene and employs a fixed questioning template to organize language to ask about which object will be manipulated. Although the use of CLIP enhances the generalization ability for object recognition, it can’t generate text for asking questions and to comprehend answers from the outside world and SeeAsk (Mo et al. (2023)) focuses solely on addressing ambiguities concerning object color and spatial relationship due to a fixed questioning template. KNOWNO (Ren et al. (2023a)) utilizes LLM to score the next action to be taken. If the score difference between the top two actions is less than a threshold, it’s considered ambiguity, prompting a confirmation for the final action. This approach improves efficiency and autonomy. Matcha (Zhao et al. (2023c)) not only employs vision but also utilizes haptic and sound senses to perceive object properties, such as material. When encountering ambiguity in object attribute recognition, it leverages LLM to generate inquiry content. CoELA (Zhang et al. (2023b)) utilizes LLM as both a communication module and a planning module to enhance interaction text generation and comprehension, as well as task scheduling, with collaborative agent. LLM-GROP (Ding et al. (2023)) utilizes LLM to extract latent commonsense knowledge embedded within task instruction. For example, a task instruction might be ”set dinner table with plate and fork,” while the latent commonsense knowledge could be ”fork is on the left of a bread plate.”
As for corrective instruction, LILAC (Cui et al. (2023)) utilizes GPT-3 to distinguish between task instruction and corrective instruction. It then employs Distil-RoBERTa to extract text features and input them into the network to modify the robot’s original trajectory. LATTE (Bucker et al. (2023)), on the other hand, employs BERT and CLIP to extract features from corrective instruction and observation images and input them into the network to modify the robot’s original trajectory. RT-H (Belkhale et al. (2024)) employs VLMs in a two-step operation, initially outputting abstract delta-pose representations like ”move left,” which are then converted into delta poses and human intervention can enable robots to adjust trajectories based on human language instruction.
@@
Figure 3: Foundation Models for Interaction Module. Interaction mainly involves the exchange between task instruction and corrective instruction. Ambiguity often arises in task instruction interaction, hence robot needs to detect ambiguities. 1) One approach is to perceive objects in a multi-modal environment and enumerate possible ambiguities based on perception information (Mo et al. (2023)). 2) Another approach involves using LLM to be the next step prediction module, which predicts and scores the next step; if the scores of the top 2 steps are less than $\delta$, it is considered that the task goal is ambiguous (Ren et al. (2023a)). 3) Strong comprehension skills are required during the transmission of corrective instruction, and the current mainstream approach involves using the encoder of LLM to extract tokens and input them into the policy to modify the original trajectory (Bucker et al. (2023)).

### v7

@@ 1 Introduction
Researchers aim to create universal robots that can seamlessly integrate into human life to boost productivity, much like those depicted in the movie ‘I, Robot’. However, a key hurdle in achieving this lies in the robots’ ability to manipulate objects in their unstructured environments according to different tasks. There is abundant literature available for improving the general manipulation capability of robots, which can be roughly categorized into model-based and learning-based approaches (zarrin2023hybrid). The real world is too diverse for universal robots and they must adapt to unstructured environments and arbitrary objects to manipulate effectively. Therefore, learning-based methods are crucial for manipulation tasks (kleeberger2020survey).
The predominant methodologies in learning-based approaches are deep learning, reinforcement learning and imitation learning. Learning-based methods have spanned from acquiring specific manipulation skills through labeled datasets like human demonstration, to acquiring abstract representations of manipulation tasks conducive to high-level planning, to exploring an object’s functionalities through interaction and encompassing various objectives in between (kroemer2021review). However, challenges persist, including 1) unnatural interaction with humans; 2) high-cost data collection; 3) limited perceptual capability; 4) non-intelligent hierarchy of skills; 5) inaccurate pre- and post-conditions & post-hoc correction; 6) unreliable skill learning; 7) poor environment transition (hu2023toward).
Foundation models are primarily pretrained on vast internet-scale datasets, enabling them to be fine-tuned for diverse tasks. Their significant advancements in vision and language processing contribute to mitigating the aforementioned challenges. Based on firoozi2023foundation and considering the different input modalities and functionalities of the models, we categorize foundation models into the following six types.
@@
    Large Language Models (LLMs) like BERT (devlin2018bert), GPT-3 (brown2020language) demonstrate the capability to generate coherent chains of thought.
@@
    Visual Foundation Models (VFMs) like SAM (kirillov2023segment) demonstrate strong segmentation capability for open-set objects.
@@
    Visual Generative Models (VGMs) like DALL-E (ramesh2021zero), Zero-1-to-3 (liu2023zero) and Sora (videoworldsimulators2024), demonstrate the capability to generate 2D images, videos or 3D meshes through text or images.
@@
    Visual-Language Models (VLMs) like GPT-4V (achiam2023gpt), CLIP (radford2021learning) showcase robust comprehension of both vision and language, such as open-set image classification and visual question answering.
@@
    Large Multimodal Models (LMMs) expand their scope beyond vision and language to create novel categories of foundation models incorporating additional modalities, such as ULIP (xue2023ulip) aligns point cloud representation to the pre-aligned image-text feature space. VLMs are a type of LMMs (firoozi2023foundation). However, due to the current literature focusing more on VLMs, this paper will treat VLMs as a separate category. To avoid confusion, LMMs in this paper refer to those that include images, language, and more modalities.
@@
    Robotic-specific Foundation Models (RFMs), like RT-X (rt_x_2023). Internet-scale dataset, such as images and text data, are suitable for pre-training visual and language models, but lack task level manipulation data. Therefore, researchers aim to train end-to-end RFMs by collecting task-level manipulation datasets to enable observations-to-action mapping.
@@
These findings underscore the potential of embedding foundation models into manipulation tasks as a viable path toward achieving general manipulation capability. However, we do not believe that a single foundation model alone can achieve general manipulation capability. Although RFMs currently represent a single-model end-to-end training approach, ensuring safety and stability, particularly in achieving an over 99% success rate in manipulation tasks, remains a challenge. Achieving over a 99% success rate in manipulation tasks is crucial, as human manipulation success rates are around 99%. Without this level of accuracy, robots can’t replace humans (nishanth2023). Therefore, drawing inspiration from the development of autonomous driving systems (hu2023planning), achieving general manipulation capability necessitates an overarching framework that encompasses multiple functional modules, with different foundation models assuming distinct roles in facilitating general manipulation capability.
@@
The ultimate general manipulation framework should be able to interact with human or other agent and control whole-body to manipulate arbitrary objects in open-world scenarios and achieve diverse manipulation tasks (mccarthy2024towards). Drawing from kroemer2021review and this general manipulation definition, we propose a comprehensive framework for general manipulation. However, the interaction between robot and human involves not only recognizing intentions but also learning new skills or improving old skills from human experts in the external world. Open-world scenarios may be static or dynamic. Objects can be either rigid or deformable. Task objectives can vary from short-term to long-term. Furthermore, tasks may necessitate different degrees of precision with respect to contact points and applied forces/torques. Although there are many challenges, achieving general manipulation can be accomplished through multiple stages. We designate the restriction of the robot’s learning capability to improving old skills and to manipulating rigid objects in static scenes in order to achieve short-horizon task objectives with low precision requirements for contact points and forces/torques as Level 0 (L0). At the same time, we believe that improving the algorithm performance of different modules in the framework can support the transition from the L0 stage to the final general manipulation. Hence, we aim to use this survey not only to enlighten scholars on the issues that foundation models can address in robot learning for manipulation but also to stimulate their exploration of the general manipulation framework and the role various foundation models can play in the general manipulation framework.
@@
di2023towards and firoozi2023foundation provide detailed descriptions of the application of foundation models in navigation and manipulation, but these lack thoughtful consideration of the relationship between foundation models across different applications. The survey most closely related to this paper is xiao2023robot. Compared to this survey, our survey focuses on the contributions of foundation models to robot learning for manipulation, proposing a comprehensive framework and detailing how foundation models can address challenges in each module of the framework.
@@ 2 Framework of Robot Learning for General Manipulation
Over the past decade, there has been a significant expansion in research concerning robot manipulation, with a focus on leveraging the growing accessibility of cost-effective robot arms and grippers to enable robots to interact directly with the environment in pursuit of their objectives. As the real world encompasses extensive variation, a robot cannot expect to possess an accurate model of its unstructured environment, the objects within it, or the skills necessary for manipulation in advance (kroemer2021review).
Early stage, robot manipulation is defined as learning a policy $\Pi$ through deep learning, reinforcement learning, or imitation learning, etc. This policy controls the robot’s joint movements and executes tasks based on observations of the environment and the robot’s state $S$, mapping to actions $\alpha$. such as Rlafford (geng2023rlafford) and Graspnet (fang2020graspnet) take point cloud as input and output the target pose. This process is represented by the Skill Execution module, as shown in Fig. 2.
In the mid-term, many tasks in robotics require a series of correct actions, which are often long-horizon tasks. For example, making a cup of tea with a robot involves multiple sequential steps such as boiling water, adding a tea bag, pouring hot water, etc. Learning to plan for long-horizon tasks is a central challenge in episodic learning problems (wang2020long). Decomposing tasks has several advantages. It makes learning individual skills more efficient by breaking them into shorter-horizon, thus aiding exploration. Reusing skills in multiple settings can speed up learning by avoiding the need to relearn elements from scratch each time. Researchers train a hierarchy model to decompose the task into a sequence of subgoals (ahn2022can), and observe pre- and post-conditions to ensure that the prerequisites and outcomes of each subgoals are met (cui2022can). These three processes are represented as the Hierarchy of Skills module $H$, the Pre-conditions Detection module $P$, and the Post-conditions Detection module $P$ in Fig. 2. However, detecting only task success with post-conditions detection is insufficient. It should also identify the reasons for task failure to help the robot self-correct and improve success rates. Therefore, we add a Post-hoc Correction module, as shown in Fig. 2.
Recently, researchers have realized that training policies require real-world interaction between the robot and environments, which inevitably increases the probability of unforeseen hazardous situations. Therefore, researchers aim to train the environment’s transition model $T$. Once the model is fitted, robot can generate samples based on it, significantly reducing the frequency of direct interaction between the robot and environments (liu2024learning). This process is represented as the Transition module $T$ in Fig. 2.
@@
Figure 2: Framework of Robot Learning for General Manipulation. The Pre-conditions Detection module $P$ perceives the environment to identify objects and the affordances objects support. The Interaction module $I$ receives instruction from a human or other agent. It uses perception information from the Pre-conditions Detection module $P$ to check for ambiguities in the instruction. If there are any ambiguities, it generates a question to clarify the instruction by asking the human or other agent. The Hierarchy of Skills module $H$ generates subgoals by using precise instruction from the Interaction module $I$ and perception information from the Pre-conditions Detection module $P$. Each subgoal is then passed to the Skill Execution module. In the Skill Execution module, Policy module $\Pi$ generates Action $\alpha$ based on the State $S$. To obtain the next state after executing the current action, State $S$ can either perceive it from the environment or use the Transition module $T$. To train the Skill Execution module, including the State module $S$, the Policy module $\Pi$ and the Transition module $T$, the Manipulation Data Generation module is required. This module provides a task-level manipulation dataset. When issues arise during execution, corrective instruction is sent to the Policy module $\Pi$ for manual adjustment. Policy module $\Pi$ modifies the current action to corrective action and saves corrective demonstration to the dataset for self-improvement of Policy module $\Pi$. After skill execution, Post-conditions Detection module $P$ determines the success of execution. If successful, proceed to the next subgoal; if not, the failure reason is conveyed to Post-hoc Correction module for self-correction.
@@
    Hierarchy of skills. This module takes as input the perception information about objects in the environment and their affordances for the task from the Pre-conditions Detection module $P$, as well as the precise instruction from the Interaction module $I$. It then produces a sequence of subgoals as output. The concept of ‘Hierarchy of skills’ often involves creating a sequence of subgoals (song2023llm). Each subgoal necessitates a skill, which may consist of one or multiple primitive actions (zhang2023bootstrap). For instance, tasks like filling the kettle with water, heating the water, and getting the tea leaves are examples of subgoals that robot needs to achieve in a specific order to fulfill the final goal as instructed.
@@
    State. The input to the State module is the current environment, objects and robot observation. States require the use of multiple sensors for perception. The output is the features of states. The states consists of robot proprioception $S_{robot}$, environment state $S_{e}$, and objects states $S_{o}$. The difference between $S_{e}$ and $S_{o}$ is analogous to the foreground and background of an image. $S_{robot}$ generally relates to the mechanical structure of the robot. Currently, there are limited studies focusing on the improvement of robot mechanical structure using foundation models, with stella2023can being one of them. However, researches in this direction are scarce and still in their initial stages.
@@
    Transition. The Transition module $T$ takes an action generated by the Policy module $P$ as input. It outputs the next state after executing this action, thus helping to reduce the interaction between the robot and the real environment. UniSim (yang2023learning) introduces the action-in-video-out framework as an observation prediction model. It takes the current action as input and produces the subsequent observation as output.
@@ 3 Human/Agent Interaction
There are two ways for human or other agent to interact with robot: 1) Providing task instruction to the robot to help it understand the task objective and complete the task independently (khan2023natural). 2) Collaborating with the human or other agent to complete tasks, sharing workspace information, and conveying corrective instruction when useful or error-correcting information is identified to optimize the robot’s current action (lynch2023interactive).
@@
In addressing instruction ambiguity and text generation and comprehension challenges, SeeAsk (mo2023towards) utilizes CLIP’s perceptual module to identify objects in the scene and employs a fixed questioning template to organize language to ask about which object will be manipulated. Although the use of CLIP enhances the generalization ability for object recognition, it can’t generate text for asking questions and to comprehend answers from the outside world and SeeAsk (mo2023towards) focuses solely on addressing ambiguities concerning object color and spatial relationship due to a fixed questioning template. KNOWNO (ren2023robots) utilizes LLM to score the next action to be taken. If the score difference between the top two actions is less than a threshold, it’s considered ambiguity, prompting a confirmation for the final action. This approach improves efficiency and autonomy. Matcha (zhao2023chat) not only employs vision but also utilizes haptic and sound senses to perceive object properties, such as material. When encountering ambiguity in object attribute recognition, it leverages LLM to generate inquiry content. CoELA (zhang2023building) utilizes LLM as both a communication module and a planning module to enhance interaction text generation and comprehension, as well as task scheduling, with collaborative agent. LLM-GROP (ding2023task) utilizes LLM to extract latent commonsense knowledge embedded within task instruction. For example, a task instruction might be ”set dinner table with plate and fork,” while the latent commonsense knowledge could be ”fork is on the left of a bread plate.”
As for corrective instruction, LILAC (cui2023no) utilizes GPT-3 to distinguish between task instruction and corrective instruction. It then employs Distil-RoBERTa to extract text features and input them into the network to modify the robot’s original trajectory. LATTE (bucker2023latte), on the other hand, employs BERT and CLIP to extract features from corrective instruction and observation images and input them into the network to modify the robot’s original trajectory. RT-H (belkhale2024rt) employs VLMs in a two-step operation, initially outputting abstract delta-pose representations like ”move left,” which are then converted into delta poses and human intervention can enable robots to adjust trajectories based on human language instruction.
@@
Figure 3: Foundation Models for Interaction Module. Interaction mainly involves the exchange between task instruction and corrective instruction. Ambiguity often arises in task instruction interaction, hence robot needs to detect ambiguities. 1) One approach is to perceive objects in a multi-modal environment and enumerate possible ambiguities based on perception information (mo2023towards). 2) Another approach involves using LLM to be the next step prediction module, which predicts and scores the next step; if the scores of the top 2 steps are less than $\delta$, it is considered that the task goal is ambiguous (ren2023robots). 3) Strong comprehension skills are required during the transmission of corrective instruction, and the current mainstream approach involves using the encoder of LLM to extract tokens and input them into the policy to modify the original trajectory (bucker2023latte).